# Mitigating the Backdoor Effect for Multi-Task Model Merging via Safety-Aware Subspace

### v1

@@ Abstract
Model merging has gained significant attention as a cost-effective approach to integrate multiple single-task fine-tuned models into a unified one that can perform well on multiple tasks. However, existing model merging techniques primarily focus on resolving conflicts between task-specific models, they often overlook potential security threats, particularly the risk of backdoor attacks in the open-source model ecosystem. In this paper, we first investigate the vulnerabilities of existing model merging methods to backdoor attacks, identifying two critical challenges: backdoor succession and backdoor transfer. To address these issues, we propose a novel Defense-Aware Merging (DAM) approach that simultaneously mitigates task interference and backdoor vulnerabilities. Specifically, DAM employs a meta-learning-based optimization method with dual masks to identify a shared and safety-aware subspace for model merging. These masks are alternately optimized: the Task-Shared mask identifies common beneficial parameters across tasks, aiming to preserve task-specific knowledge while reducing interference, while the Backdoor-Detection mask isolates potentially harmful parameters to neutralize security threats. This dual-mask design allows us to carefully balance the preservation of useful knowledge and the removal of potential vulnerabilities. Compared to existing merging methods, DAM achieves a more favorable balance between performance and security, reducing the attack success rate by 2-10 percentage points while sacrificing only about 1% in accuracy. Furthermore, DAM exhibits robust performance and broad applicability across various types of backdoor attacks and the number of compromised models involved in the merging process. We will release the codes and models soon.
@@ 1 Introduction
The rapid advancement of artificial intelligence has led to the emergence of pre-trained models that demonstrate exceptional performance across various tasks (Yang et al., 2024a). However, training and deploying individual models for each specific task not only incurs substantial computational costs but also results in knowledge redundancy and storage inefficiencies. To address these challenges, multi-task model merging, as a promising solution, integrates parameters from multiple single-task models into a unified model (Tang et al., 2024), which not only enhances task-specific performance but also significantly improves computational efficiency and cost-effectiveness (Izmailov et al., 2018; Frankle et al., 2020; Ilharco et al., 2022b).
@@ 2.1 Preliminaries
Previous Multi-Task Merging Techniques. Denote the $f_{\theta}$ as the CLIP-like pre-trained model $f$ with weights ${\theta}$ and a set of datasets $\mathcal{D}=\{D_{i}\}_{i=1}^{n}$ for $n$ downstream tasks. We can fine-tune the pre-trained model parameterized by $\theta_{\text{pre}}$ to acquire $n$ task-specific models parameterized by $\{\theta_{i}\}_{i=1}^{n}$. Then, for each task $i$, the task vector can defined as the difference between $\theta_{\text{pre}}$ and $\theta_{i}$, i.e., $\tau_{i}=\theta_{i}-\theta_{\text{pre}}$. Existing merging methods can be formulated as the optimization to acquire $\theta_{merged}=\theta_{\text{pre}}+\sum_{i=1}^{n}\{\lambda_{i}{{\tau_{i}}^{\prime}}\}$, where ${\forall}\lambda\in[0,1]$ refers to the merging coefficient and $\phi(\tau_{i})={\tau_{i}}^{\prime}$ represents the revision for each task vector. The main difference among these methods exists in ways to acquire the ${\tau_{i}}^{\prime}$ and $\lambda_{i}$. For example, both Weight Average (Wortsman et al., 2022) and Task Arithmetic (Ilharco et al., 2022a) adopt the origin task vector $\tau_{i}$, with the $\lambda=\frac{1}{n}$ adapted to the number of tasks and a fixed $\lambda=0.3$ respectively. Ties-Merging (Yadav et al., 2024) and Concrete (Tang et al., 2023) address the interference among tasks and replace the original task vector with ${\tau_{i}}^{\prime}$. Moreover, RegMean (Jin et al., 2022) and AdaMerging (Yang et al., 2023) respectively formulate the optimization of $\lambda_{i}$ according to the model’s activations and the entropy on an unlabeled held-out dataset. However, these works share the same and single optimization objective to maximize the performance of the merging model on the clean test datasets as Eq.1, where $\mathcal{A}$ is a model merging algorithm associated with $\phi(\cdot)$ and $\lambda$. It is uncertain if current merging methods remain effective considering safety issues like backdoors, which introduce new evaluation requirements for merging algorithms.
@@
Merging Considering the Existence of Backdoor. Typically, a model injected with the backdoor behaves normally for clean input data but will be misguided toward the target class when the inputs contain a specific trigger (Wu et al., 2024). Thus, when backdoored task-specific models are utilized during merging, the Eq.1 should also be rewritten as Eq.2. Universally, the performance on the clean test dataset and the safety evaluated on the test dataset with the trigger can be respectively defined as the accuracy (ACC) and attack success rate (ASR) (Wu et al., 2024). In other words, the optimization of model merging should be towards high ACC and low ASR simultaneously to maximize safety and performance considering the backdoor.
$\max_{\phi,\lambda}\frac{1}{n}\sum_{i=1}^{n}\left(\text{Performance}\left(\mathcal{A}(\theta_{\text{pre}},\phi(\tau_{i}),\lambda_{i}),D_{i}^{\text{test}}\right)+\text{Safety}\left(\mathcal{A}(\theta_{\text{pre}},\phi(\tau_{i}),\lambda_{i}),D_{i}^{\text{test with trigger}}\right)\right)$ (2)
@@ 2.2 Unpacking the Phenomenon of Backdoor Succession and Backdoor Transfer
Based on Eq.2, we further explore whether existing merging methods can naturally cope with neglected backdoor issues. Specifically, we take the CLIP-ViT (Radford et al., 2021) as the pre-trained model and explore the backdoor effect during model merging on image classification tasks (Tang et al., 2024).
Backdoored Model Constructions: We utilize the commonly used vit-specific backdoor attack, TrojVit (Zheng et al., 2023), to construct backdoored models. As shown in Figure 1 and Figure 6, backdoored models achieve good ACC but higher (worse) ASR than clean ones, bringing the potential risk to model merging. We present the ASR for clean and backdoored models to demonstrate the backdoor impact (Zheng et al., 2023).
The original object of multi-task merging is to provide a cost-effective parameter-level fusion strategy to obtain a merged multi-task model that can achieve close or better performance than individual fine-tuned models for each task (Yang et al., 2024b). While considering the existence of the backdoor during merging, this object can be transformed as Promote the merged model close or even upper to the ACC of clean individual fine-tuned models and lower or at least equal to the ASR of the backdoored individual fine-tuned models. Thus, we explore the backdoor effect in multi-task merging by comparing the ACC and ASR of the merged model with those of individual fine-tuned models across tasks.
Backdoor Succession During Merging: As shown in Figure 2 and Figure 7, we observe that the development of merging techniques has increased the ACC of merged model close or even better than individual fine-tuned models. But unfortunately, from the perspective of safety evaluation(ASR), we have the Finding 1.
@@
Backdoor Transfer During Merging: Take Figure 3(b) for an example, we provide the clean task-specific model for merging, but the merged models adopting existing merging methods all achieve worse(higher) ASR than the provided clean task-specific model on the relative task. This can be accounted for by Finding 2.
Finding 2: Backdoor Transfer: The backdoor effect from backdoored task-specific models can transfer to other clean task-specific models during multi-task merging.
@@ 3 Defense-Aware Merging
Based on the analysis shown in Section 2, we aim to weaken the backdoor effect and provide a defense-ware merging (DAM) algorithm to achieve a merged model with high ACC and low ASR. To achieve this goal, we face two challenges: (i) How can we cope with backdoor issues during merging when we can not know the backdoor types and whether the models that need merging are safe or not in advance? (ii) How can we unify the optimization process to achieve a good trade-off between performance and safety during merging?
To address (i), we respectively synthesize a universal perturbation for each task-specific model to represent undesired behavioral changes from trigger insertion, without requiring assumptions about backdoor information (e.g. the trigger’s size or location) (Zeng et al., 2024). Assisted by the synthesized perturbations, we can identify and adjust the parameters related to the backdoor during merging, assuming that the backdoor-related parameters are more sensitive to the perturbations (Wu et al., 2024). To address (ii), we provide a dual-mask optimization strategy to identify a shared and safety-aware subspace on task vectors, which can concurrently mitigate interference and backdoor issues during merging with the safe and unsafe components discerned using the learned perturbations from (i), achieving a good trade-off between performance and safety.
Thus, the framework of DAM can be formulated as a bi-level optimization problem as Eq. 3:
@@
where $\mathcal{A}$ is the merging operation associated with $\lambda$ and two different mask designs (Task-Shared Mask $M_{1}$ and Backdoor-Detection Mask $M_{2}$), where $M_{1},M_{2}\in\mathbb{R}^{d}$ referring to the $\phi(\cdot)$ in Eq 1 to revise the task vectors $\tau$. The $M_{1}\odot M_{2}$ aims to achieve a union set for two masks through the element multiplication. The $\alpha$ is a balance weight for mask optimization, with larger $\alpha$ values favoring safety over performance. The $\mathcal{L}$ is usually defined as the unsupervised loss such as entropy loss as Eq. 4 for test-time adaptation under the share-and-play scenario (Yang et al., 2023) where we have no access to the training data. For image classification tasks, we can initially employ the merged model to generate predictions $\hat{y}$ on the unlabeled test data and subsequently utilize these predictions to optimize the merged model, where $x_{i}$ is the $i$-th unlabeled sample, $p(\hat{y}_{c}|x_{i})$ is the predicted probability of the $c$-th class, and $C$ is the number of classes.
@@
We can verify the reasonability of this dual-mask strategy from two aspects: (i) From the perspective of Model Merging, the design of $M_{1}$ assumes that the interference among tasks is the key to influencing merged model performance and it usually exists in the shared parameter space among different individual finetuned models (Tang et al., 2023). As shown in Figure 4: The $M_{1}$ is designed to identify the shared parameter space among different task vectors. Then, the task-specific parameters can be separated after Subspace Masking I. Following the purple Merging Flow, the parameter of the final merged model includes two parts: the merged version of these task-specific parameters and the parameter from the pre-trained model corresponding to the shared mask $M_{1}$ region. However, though it has been verified that revising the task vectors through mask designs similar to $M_{1}$ can improve the merged model performance clearly (Tang et al., 2024), the backdoor will still succeed as shown in the outcome of the purple Merging Flow. That’s because the backdoor-related parameters (the red cross) may exist in the task-specific parameter space (blue part). Thus, we propose to utilize an additional mask $M_{2}$ to identify these backdoor-related parameters and then integrate with $M_{1}$ to acquire the shared and safety-aware subspace. After the Subspace Masking II, following the yellow Merging Flow, the backdoor effect from fine-tuned models can be mitigated by replacing the backdoor-related parameters with pre-trained weights. But notably, the introduced $M_{2}$ will also lose part of the useful task-specific parameter (the green and yellow part existing in the represented region of $M_{2}$), leading to a decrease in model performance. Thus, the alternate optimization of two masks can be seen as seeking a good trade-off between performance and safety during merging; (ii) From the perspective of Pareto Optimal Balance between performance and safety during the dual mask optimization, we propose Theorem 1.
@@
Proof of Theorem 1.
Moreover, we provide the convergence analysis to discuss how the proposed algorithm converges to a Pareto optimal solution to balance the performance and safety in the Appendix B.3.
@@ 4.1 Experimental Setup
Datasets and Models: Following  (Tang et al., 2024), we utilize CLIP-ViT-B/32 and CLIP-ViT-L/14 as our pre-trained models and conduct experiments on six image classification tasks including Stanford Cars (Krause et al., 2013), RESISC45 (Cheng et al., 2017), EuroSAT (Helber et al., 2018), SVHN (Netzer et al., 2021), MNIST (Lecun et al., 1998) and DTD (Cimpoi et al., 2014). We first construct the clean individual fine-tuned models by directly fine-tuning the pre-trained model on these clean datasets and then inject them with the backdoor adopting TrojVit (Zheng et al., 2023) and BadVit (Yuan et al., 2023) strategy to construct the backdoored models. More detailed descriptions of datasets and models can be shown in the Appendix C.1.
Baselines: We mainly compare DAM with three groups of baselines:(i) Individual Finetuning: All Clean Finetuned Models, All Backdoored Finetuned Models, and Mixing with Clean and Backdoored Finetuned Models under different settings. Notably, we just average their results for reference;(ii) Multi-Task Merging Methods: Weight Average (Wortsman et al., 2022), Fisher merging (Matena & Raffel, 2022), RegMean (Jin et al., 2022), Task Arithmetic (Ilharco et al., 2022a), Ties-Merging (Yadav et al., 2024), Adamerging (Yang et al., 2023), Concrete (Tang et al., 2023) (iii) Post-Defense Methods involving adversarial perturbation: ANP (Wu & Wang, 2021), AWM (Chai & Chen, 2022) and SAU (Wei et al., 2023)). Notably, we execute these post-defense backdoor processing on the best-merged model adopting (ii), which can be seen as two-stage methods compared with our end-to-end training process. Moreover, we also provide experiments and discussions with other backdoor-related merging works (e.g.LoRA-as-an-Attack (Liu et al., 2024) and BadMerging (Zhang et al., 2024).), which are not designed for multi-task merging to further clarify our contributions. More detailed discussions can be shown in Appendix A.
@@
Evaluation Metric: We respectively adopt the top-1 accuracy(ACC) on clean test data as a performance metric and the attack success rate(ASR) on test data with trigger as a safety metric  (Wu et al., 2024). An ideal model should have high ACC but low ASR. To further explore the backdoor effect, apart from the average ACC and ASR on the full six tasks, we also present the average results on tasks related to backdoored task-specific models, including ACC(2)/ASR(2) and ACC(4)/ASR(4), with the numbers indicating the count of backdoored models.

### v2

@@ Abstract
Model merging has gained significant attention as a cost-effective approach to integrate multiple single-task fine-tuned models into a unified one that can perform well on multiple tasks. However, existing model merging techniques primarily focus on resolving conflicts between task-specific models, they often overlook potential security threats, particularly the risk of backdoor attacks in the open-source model ecosystem. In this paper, we first investigate the vulnerabilities of existing model merging methods to backdoor attacks, identifying two critical challenges: backdoor succession and backdoor transfer. To address these issues, we propose a novel Defense-Aware Merging (DAM) approach that simultaneously mitigates task interference and backdoor vulnerabilities. Specifically, DAM employs a meta-learning-based optimization method with dual masks to identify a shared and safety-aware subspace for model merging. These masks are alternately optimized: the Task-Shared mask identifies common beneficial parameters across tasks, aiming to preserve task-specific knowledge while reducing interference, while the Backdoor-Detection mask isolates potentially harmful parameters to neutralize security threats. This dual-mask design allows us to carefully balance the preservation of useful knowledge and the removal of potential vulnerabilities. Compared to existing merging methods, DAM achieves a more favorable balance between performance and security, reducing the attack success rate by 2-10 percentage points while sacrificing only about 1% in accuracy. Furthermore, DAM exhibits robust performance and broad applicability across various types of backdoor attacks and the number of compromised models involved in the merging process. Our codes and models can be accessed through DAM.
@@ 1 Introduction
The rapid advancement of artificial intelligence has led to the emergence of pre-trained models that demonstrate exceptional performance across various tasks (Yang et al., 2024a). However, training and deploying individual models for each specific task not only incurs substantial computational costs but also results in knowledge redundancy and storage inefficiencies. To address these challenges, multi-task model merging, as a promising solution, integrates parameters from multiple single-task models into a unified model (Tang et al., 2024a), which not only enhances task-specific performance but also significantly improves computational efficiency and cost-effectiveness (Izmailov et al., 2018; Frankle et al., 2020; Ilharco et al., 2022b).
@@ 2.1 Preliminaries
Previous Multi-Task Merging Techniques. Denote the $f_{\theta}$ as the CLIP-like pre-trained model $f$ with weights ${\theta}$ and a set of datasets $\mathcal{D}=\{D_{i}\}_{i=1}^{n}$ for $n$ downstream tasks. We can fine-tune the pre-trained model parameterized by $\theta_{\text{pre}}$ to acquire $n$ task-specific models parameterized by $\{\theta_{i}\}_{i=1}^{n}$. Then, for each task $i$, the task vector can defined as the difference between $\theta_{\text{pre}}$ and $\theta_{i}$, i.e., $\tau_{i}=\theta_{i}-\theta_{\text{pre}}$. Existing merging methods can be formulated as the optimization to acquire $\theta_{merged}=\theta_{\text{pre}}+\sum_{i=1}^{n}\{\lambda_{i}{{\tau_{i}}^{\prime}}\}$, where ${\forall}\lambda\in[0,1]$ refers to the merging coefficient and $\phi(\tau_{i})={\tau_{i}}^{\prime}$ represents the revision for each task vector. The main difference among these methods exists in ways to acquire the ${\tau_{i}}^{\prime}$ and $\lambda_{i}$. For example, both Weight Average (Wortsman et al., 2022) and Task Arithmetic (Ilharco et al., 2022a) adopt the origin task vector $\tau_{i}$, with the $\lambda=\frac{1}{n}$ adapted to the number of tasks and a fixed $\lambda=0.3$ respectively. Ties-Merging (Yadav et al., 2024) and Concrete (Tang et al., 2023) address the interference among tasks and replace the original task vector with ${\tau_{i}}^{\prime}$. Moreover, RegMean (Jin et al., 2022) and AdaMerging (Yang et al., 2023) respectively formulate the optimization of $\lambda_{i}$ according to the model’s activations and the entropy on an unlabeled held-out dataset. However, these works share the same and single optimization objective to maximize the performance of the merged model on the clean test datasets as Eq.1 from the evaluation perspective, where $\mathcal{A}$ is a model merging algorithm associated with $\phi(\cdot)$ and $\lambda$. It is uncertain if current merging methods remain effective considering safety issues like backdoors, which introduces potential but important concerns for deploying merging algorithms to more scenarios.
@@
Merging Considering the Existence of Backdoor. Typically, a model injected with the backdoor behaves normally for clean input data but will be misguided toward the target class when the inputs contain a specific trigger (Wu et al., 2022). Thus, when backdoored task-specific models are utilized during merging, the Eq.1 should also be rewritten as Eq.2. Universally, the accuracy (ACC) represents the percentage of test input images without triggers classified into their corresponding correct classes (true label), while the attack success rate (ASR) shows the percentage of input images embedded with a trigger classified into the pre-defined target class (label decided by the attackers) (Wu et al., 2022). Ideally, the optimization of model merging should be towards high ACC and low ASR simultaneously to maximize safety and performance considering the existence of the backdoor. The $\omega$ is set as the balance weight for the performance and safety by default.
$\max_{\phi,\lambda}\frac{1}{n}\sum_{i=1}^{n}\left(\text{Performance}\left(\mathcal{A}(\theta_{\text{pre}},\phi(\tau_{i}),\lambda_{i}),D_{i}^{\text{test}}\right)+\omega\cdot\text{Safety}\left(\mathcal{A}(\theta_{\text{pre}},\phi(\tau_{i}),\lambda_{i}),D_{i}^{\text{test with trigger}}\right)\right)$ (2)
@@
2.2 The Settings of Model Merging Considering the Backdoor
Based on Eq.2, we further explore whether existing merging methods can naturally cope with neglected backdoor issues. Specifically, we take the CLIP-ViT (Radford et al., 2021) as the pre-trained model and explore the backdoor effect during model merging on image classification tasks (Tang et al., 2024a).
Backdoored Model Constructions: We utilize the commonly used vit-specific (patch-wise) backdoor attack, TrojVit (Zheng et al., 2023), to construct backdoored models. As shown in Figure 1 and Figure 6, backdoored models achieve good ACC but higher (worse) ASR than clean ones, bringing the potential risk to model merging. The detailed construction implementations can be shown in Appendix C.1.
Attack and Defense Scenario: We assume that the adversary is the provider of backdoored models and doesn’t know other task-specific models and merging algorithms. For defenders, they only have different checkpoints without the knowledge of whether or not and which region they are injected with a backdoor.
@@ 2.3 Unpacking the Phenomenon of Backdoor Succession and Backdoor Transfer
The original object of multi-task merging is to provide a cost-effective parameter-level fusion strategy to obtain a multi-task model that can achieve close or better performance than individual fine-tuned models for each task (Yang et al., 2024b). While considering the existence of the backdoor during merging, this object can be transformed as Promote the merged model close or even upper to the ACC of clean individual fine-tuned models and lower or at least equal to the ASR of the backdoored individual fine-tuned models. Thus, we explore the backdoor effect in multi-task merging by comparing the ACC and ASR of the merged model with those of individual fine-tuned models across tasks. We have the following two important findings.
Backdoor Succession During Merging: From the Figure 2 and Figure 7, we can observe that the development of merging techniques has increased the ACC of merged model close or even better than individual fine-tuned models. While considering the safety issues, the ASR of the merged model on the backdoor-related task (e.g. EUROSAT, shown in Table 14 ) decreases but is still high, due to the backdoor succession as Finding 1.
@@
Backdoor Transfer During Merging: The Figure 3 and Figure 8 describe that though we provide clean task-specific models for merging (e.g. SVHN shown in Table 14), the ASR of the merged model on SVHN will unexpectedly increase compared with the individual fine-tuned model on SVHN, due to other backdoored task-specific models. This can be accounted for by Finding 2. Specifically, from the perspective of parameter disentanglement, different task-specific models have a common parameter region, if this task-general region is injected with the backdoor, the backdoor effect can be seen as transferring from backdoored models to clean models during model merging. More detailed discussions can be shown in the Appendix B.5.
Finding 2: Backdoor Transfer: The backdoor effect from backdoored task-specific models can transfer to other clean task-specific models, which leads to special safety-related challenges for multi-task merging.
@@ 3 Defense-Aware Merging
Based on the analysis shown in Section 2, we face two challenges to achieve a merged model with high ACC and low ASR: (i) How can we cope with backdoor issues during merging when we can not know the backdoor types and whether the models that need merging are safe or not in advance? (ii) How can we unify the optimization process to achieve a good trade-off between performance and safety during merging?
To address (i), we respectively synthesize a universal perturbation for each task-specific model to represent undesired behavioral changes from trigger insertion, without requiring assumptions about backdoor information (e.g. the trigger’s size or location) (Zeng et al., 2024). Assisted by the synthesized perturbations, we can identify and adjust the parameters related to the backdoor during merging, assuming that the backdoor-related parameters are more sensitive to the perturbations (Wu et al., 2022). To address (ii), we provide a dual-mask optimization strategy to identify a shared and safety-aware subspace on task vectors, which represents the low-dimensional-parameter area compared with the full parameters of task vectors, aiming to concurrently mitigate interference and backdoor issues during merging with the safe and unsafe components discerned using the learned perturbations from (i), achieving a good trade-off between performance and safety.
Thus, the framework of DAM can be formulated as a bi-level optimization problem as Eq. 3, where $\mathcal{A}$ is the merging operation associated with $\lambda$ and two different mask designs (Task-Shared Mask $M_{1}$ and Backdoor-Detection Mask $M_{2}$), where $M_{1},M_{2}\in\mathbb{R}^{d}$ referring to the $\phi(\cdot)$ in Eq 1 to revise the task vectors $\tau$. The $M_{1}\odot M_{2}$ aims to achieve a union set for two masks through the element multiplication. The $\alpha$ is a balance weight for mask optimization, with larger $\alpha$ values favoring safety over performance.
@@
The $\mathcal{L}$ is usually defined as the unsupervised loss such as entropy loss as Eq. 4 for test-time adaptation under the share-and-play scenario (Yang et al., 2023) where we have no access to the training data. For image classification tasks, we can initially employ the merged model to generate predictions $\hat{y}$ on the unlabeled test data and subsequently utilize these predictions to optimize the merged model, where $x_{i}$ is the $i$-th unlabeled sample, $p(\hat{y}_{c}|x_{i})$ is the predicted probability of the $c$-th class, and $C$ is the number of classes.
@@
We can verify the reasonability of this dual-mask strategy from two aspects: (i) From the perspective of Model Merging, the design of $M_{1}$ assumes that the interference among tasks is the key to influencing merged model performance and it usually exists in the shared parameter space among different individual finetuned models (Tang et al., 2023). As shown in Figure 4: The $M_{1}$ is designed to identify the shared parameter space among different task vectors. Then, the task-specific parameters can be separated after Subspace Masking I. Following the purple Merging Flow, the parameter of the final merged model includes two parts: the merged version of these task-specific parameters and the parameter from the pre-trained model corresponding to the shared mask $M_{1}$ region. However, though it has been verified that revising the task vectors through mask designs similar to $M_{1}$ can improve the merged model performance clearly (Tang et al., 2024a), the backdoor will still succeed as shown in the outcome of the purple Merging Flow. That’s because the backdoor-related parameters (the red cross) may exist in the task-specific parameter space (blue part). Thus, we propose to utilize an additional mask $M_{2}$ to identify these backdoor-related parameters and then integrate with $M_{1}$ to acquire the shared and safety-aware subspace. After the Subspace Masking II, following the yellow Merging Flow, the backdoor effect from fine-tuned models can be mitigated by replacing the backdoor-related parameters with pre-trained weights. But notably, the introduced $M_{2}$ will also lose part of the useful task-specific parameter (the green and yellow part existing in the represented region of $M_{2}$), leading to a decrease in model performance. Thus, the alternate optimization of two masks can be seen as seeking a good trade-off between performance and safety during merging; (ii) From the perspective of Pareto Optimal Balance between performance and safety during the dual mask optimization, we propose Theorem 1.
@@
The detailed proof can be shown in Appendix B.3. Moreover, we also provide the convergence analysis to discuss how DAM converges to a Pareto optimal solution to balance performance and safety.
@@ 4.1 Experimental Setup
Datasets and Models: Following  (Tang et al., 2024a), we utilize CLIP-ViT-B/32 and CLIP-ViT-L/14 as our pre-trained models and conduct experiments on six image classification tasks including Stanford Cars (Krause et al., 2013), RESISC45 (Cheng et al., 2017), EuroSAT (Helber et al., 2018), SVHN (Netzer et al., 2021), MNIST (Lecun et al., 1998) and DTD (Cimpoi et al., 2014). We first construct the clean individual fine-tuned models by directly fine-tuning the pre-trained model on these clean datasets and then inject them with the backdoor adopting TrojVit (Zheng et al., 2023) and BadVit (Yuan et al., 2023) strategy to construct the backdoored models. More detailed descriptions of datasets and models can be shown in the Appendix C.1.
Baselines::(i) Individual Finetuning: All Clean Finetuned Models, All Backdoored Finetuned Models, and Mixing with Clean and Backdoored Finetuned Models under different settings. Notably, we just average their results for reference;(ii) Multi-Task Merging Methods: Weight Average (Wortsman et al., 2022), Fisher merging (Matena & Raffel, 2022), RegMean (Jin et al., 2022), Task Arithmetic (Ilharco et al., 2022a), Ties-Merging (Yadav et al., 2024), Adamerging (Yang et al., 2023), Concrete (Tang et al., 2023) (iii) Post-Defense Methods involving adversarial perturbation: ANP (Wu & Wang, 2021), AWM (Chai & Chen, 2022) and SAU (Wei et al., 2023)). Notably, we execute these post-defense backdoor processing on the best-merged model adopting (ii), which can be seen as two-stage methods compared with our end-to-end training process. (iv) Other backdoor-related merging works that are not designed for multi-task merging: Both WAG (Arora et al., 2024) and LoRA-as-an-Attack (Liu et al., 2024) defend the backdoor by directly averaging the homogeneous clean and backdoored full model weights or LoRa on the same task without other complex designs (e.g.subspace). Moreover, the BadMerging (Zhang et al., 2024) is a newly proposed backdoor attack adapted to model merging. More detailed discussions can be shown in Appendix A.
@@
Evaluation Metric: We respectively adopt the top-1 accuracy(ACC) on clean test data as a performance metric and the attack success rate(ASR) on test data with trigger as a safety metric  (Wu et al., 2022). An ideal model should have high ACC but low ASR. To further explore the backdoor effect, apart from the average ACC and ASR on the full six tasks, we also present the average results on tasks related to backdoored task-specific models, including ACC(2)/ASR(2) and ACC(4)/ASR(4), with the numbers indicating the count of backdoored models.